OpenStudy (anonymous):

how do you find the geometric sum of (2/5)^(2k)

OpenStudy (campbell_st):

let k = 1... 1st term = (2/5)^2 2nd term = (2/5)^4 3rd term (2/5)^6 r = 4/25 since |r| < 1 then you can find the limiting sum using \[\S = a/(1 - r)\] S = (4/25)/(1 - 4/25) S = 4/25 x 25/21 The limiting sum is 4/21

OpenStudy (anonymous):

thank you! i get it